Calcium phosphate, silicon dioxide, and coke may be heated together in an electric furnace to produce phosphorus, as shown in the following equation:
2 Ca3(PO4)2 + 6 SiO2 + 10C--> 6CaSiO3 + 10 CO + P4
In this reaction, how many pounds of calcium phosphate would be needed to make 100 lb of phosphorus? I got 500
How much calcium silicate would be formed as a by product? How much sand would be used up in this same reaction?
Answer #1

100 lb = 45.3592 Kg of Phosphours P4

Molar mass of P4 = 123.90 gm/mol

Number of moles of P4 = 45359.2/123.90 = 366.09 moles

Moles of Ca3(PO4)2 required = 2 * moles of P4 = 2 * 366.09 = 732.19 moles

Molar mass of Ca3(PO4)2 = 310.18 gm/mol

Mass of Ca3(PO4)2 required = 732.19 * 310.18 gm = 227110.6942 gms = 227.110 Kg

Mass in Lbs = 500.69 Lbs

number of moles of CaSiO3 formed = 366.09 * 6 = 2196.54 moles

Molar mass of CaSiO3 = 124.33 gm/mol

Weight of CaSiO3 formed = 2196.54 * 124.33 = 273095.81 gms = 273.095 Kg = 602.07 pounds

number of moles of SiO2 required= 366.09 * 6 = 2196.54 moles

Molar mass of SiO2 = 60.08 gm/mol

Weight of SiO2 required = 2196.54 * 60.08 = 131968.12 gms = 131.968 Kg =290.939 pounds